OpenStudy (anonymous):

The slopes of three lines are given below. m = -1/2 m = 3 m = 0 a. Determine the slope of a line that is parallel to a line with each given slope. b. Determine the slope of a line that is perpendicular to a line with each given slope.

OpenStudy (anonymous):

@Littlebird @DanJS @praetorian.10

OpenStudy (danjs):

A parallel line has the same slope as the other line that it is parallel to

OpenStudy (danjs):

The slope of a perpendicular line will have a slope that is the negative reciprical of the given lines slope... IE ---> slope = 5 Perpendicular slope = -1/5 change the sign and flip the fraction over

OpenStudy (anonymous):

So for m = -1/2 Parallel: m = -1/2 Perp: m = 2

OpenStudy (littlebird):

Important note: A line perpendicular with the slope of 0 is undefined.

OpenStudy (littlebird):

I need to reword that

OpenStudy (anonymous):

No I get it!

OpenStudy (anonymous):

I remember my teacher going over undefined slopes

OpenStudy (danjs):

yeah you hve the first right

OpenStudy (anonymous):

m = 3 Parallel: 3 Perp: -1/3

OpenStudy (danjs):

yep

OpenStudy (anonymous):

Do I have to write an actual like, like y = -1/3 + 2 or do I just write the slopes

OpenStudy (anonymous):

It says determine the slope of a line for them, so I'm thinking just the slope

OpenStudy (danjs):

It is just asking for the slopes, you would need a point on the line to write its equation.

OpenStudy (anonymous):

Thanks for the clarification @DanJS As for the last one of m = 0 Parallel: 0 Perp: Undefined

OpenStudy (danjs):

right, zero slope is horizontal line, if you have a line perpendicular to that , it is a vertical line. the slope of a vertical line is infinity, or undefined

OpenStudy (anonymous):

Couldn't have found the perp for that one without your advice @Littlebird

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

Thanks both of you @Littlebird @DanJS